OpenStudy (anonymous):

Which expression is equivalent to (82)−4? 1 over 8 to the 2nd negative 1 over 8 times 8 times 8 times 8 times 8 times 8 times 8 times 8 1 over 8 times 8 times 8 times 8 times 8 times 8 times 8 times 8 82

OpenStudy (anonymous):

its not 82 its 8 squared

OpenStudy (anonymous):

raised to the -4th

OpenStudy (anonymous):

@MikeyMaximum

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

@texaschic101 i really need help

OpenStudy (anonymous):

@barreraA

OpenStudy (anonymous):

anyone?

OpenStudy (texaschic101):

(8^2) ^ -4 I am not too sure if this is right... 1/ (8^2) ^ 4 = 1/ (8^8)

OpenStudy (anonymous):

so how do i do it?

OpenStudy (anonymous):

can you just tell me the answer pls?

OpenStudy (anonymous):

like included in the how to solve it?

OpenStudy (texaschic101):

I am not even sure if it is right...not too good with exponents (8^2) ^ -4 ( I used this rule...a^-n = 1/a^n....so that turned it into this: 1/(8^2) ^ 4.....(I used this rule for this step....(a^m)^n = a^mn...and got this : 1/8^8......or 1/(8 * 8 * 8 * 8 * 8 * 8 * 8 * 8)

OpenStudy (anonymous):

lol ok thanks sooo much lol ur the only person who has actually helped

OpenStudy (texaschic101):

I try :)

OpenStudy (anonymous):

i got it right!!

OpenStudy (anonymous):

thx bye